# Mipha: A Comprehensive Overhaul of Multimodal Assistant with Small Language Models

### 15 Mar 2024 v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have showcased impressive skills in tasks related to visual understanding and reasoning. Yet, their widespread application faces obstacles due to the high computational demands during both the training and inference phases, restricting their use to a limited audience within the research and user communities. In this paper, we investigate the design aspects of Multimodal Small Language Models (MSLMs) and propose an efficient multimodal assistant named Mipha, which is designed to create synergy among various aspects: visual representation, language models, and optimization strategies. We show that without increasing the volume of training data, our Mipha-3B outperforms the state-of-the-art large MLLMs, especially LLaVA-1.5-13B, on multiple benchmarks. Through detailed discussion, we provide insights and guidelines for developing strong MSLMs that rival the capabilities of MLLMs. Our code is available at https://github.com/zhuyiche/Mipha.
@@ 4.3 Optimization
Table 3: Multi-modal evaluation on academic-task-oriented benchmarks. VQAv2 Goyal et al. (2017); GQA Hudson and Manning (2019); VizWiz Gurari et al. (2018); $\text{SQA}^{\text{I}}$: ScienceQA-IMG Lu et al. (2022); $\text{VQA}^{\text{T}}$: TextVQA Singh et al. (2019). $*{}^{*}$he training images/annotations of the datasets are observed during training. ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Column Res. is the image resolution of visual backbone. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
[…]
Multimodal Large Language Models
BLIP-2 Li et al. (2023c) Vicuna (13B) 224 129M - 65.0 41.0 19.6 61.0 42.5
InstructBLIP Dai et al. (2023b) Vicuna (7B) 224 129M 1.2M - 49.2 34.5 60.5 50.1
InstructBLIP Dai et al. (2023b) Vicuna (13B) 224 129M 1.2M - 49.5 33.4 63.1 50.7
Shikra Chen et al. (2023b) Vicuna (13B) 224 600K 5.5M $*{}^{*}$
IDEFICS-9B Laurençon et al. (2023) LLaMA (7B) 224 353M 1M 50.9 38.4 35.5 - 25.9
IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 36.0 - 30.9
Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 54.5 68.7 58.2
LLaVA-1.5 Liu et al. (2023a) Vicuna (7B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 50.0 66.8 58.2
LLaVA-1.5 Liu et al. (2023a) Vicuna (13B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 53.6 71.6 61.3
Multimodal Small Language Models
MobileVLM Chu et al. (2023) M-LLaMA (2.7B) 336 558K 665K - 59.0 - 61.2 47.5
MobileVLM-v2 Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 - 70.0 57.5
MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 43.7 70.3 57.0
Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 - 70.9 -
Mipha-3B (Ours) Phi-2 (2.7B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 47.5 70.9 56.6
@@ 5.2 Comparison with SOTA Models
Table 4: Multi-modal evaluation on instruction following benchmarks. POPE Li et al. (2023d); MME-P: MME Perception Fu et al. (2023); MME-C: MME Cognition Fu et al. (2023); MMB: MMBench Liu et al. (2023d); SEED: SEED-Bench Li et al. (2023b); MM-Vet Yu et al. (2023). ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
Method PT IT POPE MME- MMB SEED MM-Vet
P C en cn img
Multimodal Large Language Models
BLIP2-14B Li et al. (2023c) 129M - 85.3 1293.8 290.0 - - 49.7 22.4
InstructBLIP-8B Dai et al. (2023b) 129M 1.2M - - - 36 23.7 58.8 26.2
InstructBLIP-14B Dai et al. (2023b) 129M 1.2M 78.9 1212.8 291.8 - - - 25.6
Shikra-13B Chen et al. (2023b) 600K 5.5M - - - 58.8 - - -
IDEFICS-80B Laurençon et al. (2023) 353M 1M - - - 54.5 38.1 53.2 -
Qwen-VL Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
Qwen-VL-Chat Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
mPLUG-Owl2-7B Ye et al. (2023b) 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
LLaVA-7B Liu et al. (2023b) 595K 158K 72.9 809.6 247.9 38.7 36.4 37.0 25.5
LLaVA-1.5-7B Liu et al. (2023a) 558K 665K 85.9 1510.7 316.1 64.3 58.3 66.1 30.5
LLaVA-1.5-13B Liu et al. (2023a) 558K 665K 85.9 1531.3 295.4 67.7 63.6 68.2 35.4
Multimodal Small Language Models
MobileVLM-3B Chu et al. (2023) 558K 665K 84.9 1288.9 - 59.6 - - -
MobileVLM-v2-3B Chu et al. (2024) 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
MoE-LLaVA-3.6B Lin et al. (2024) 558K 1.59M 85.7 1431.3 - 68.0 - - 35.9
Bunny-3B He et al. (2024) 2M 695K 86.6 1488.8 289.3 68.6 - - -
Mipha-3B (Ours) 558K 665K 86.7 1488.9 295.0 69.7 64.3 68.9 32.1
@@
Experimental Results on Instruction-Following Benchmarks. Furthermore, we evaluate our proposed Mipha on more challenging benchmarks that are designed for evaluation of MLLMs, i.e., POPE Li et al. (2023d), MME Fu et al. (2023), MMBench Liu et al. (2023d). MMBench-CN Liu et al. (2023d), SEED-Bench-img Li et al. (2023b), and MM-Vet Yu et al. (2023). As shown in Table 3, our proposed model, Mipha-3B, exceeds LLaVA-1.5-13B in performance on MMBench and MMBench-CN by 2% and 0.7%, respectively, despite both models being trained on the same dataset. Additionally, Mipha-3B demonstrates stronger resistance to hallucinations, surpassing LLaVA-1.5-13B on POPE by 0.8%. In comparison with other MSLMs, Bunny-3B, MoE-LLaVA-3.6B Lin et al. (2024), and MobileVLM-v2-3B Chu et al. (2024), we are astonished to find that despite Mipha-3B utilizing the least amount of training data, it achieved the best performance in 5 out of 6 benchmarks. In particular, on MMBench, Mipha-3B outperformed Bunny-3B by 1.1% and MoE-LLaVA by 1.7%, which are combined with the same visual backbone SigLIP and LM Phi-2-2.7B respectively. These results validate the effectiveness of our optimization strategies.
@@ 6 Conclusion
In this paper, we delve deeply into three aspects of Multimodal Small Language Models (MSLMs): the language models, visual representations, and optimization strategy. On the basis of our thorough analysis, we identify several improvements that can enhance the performance of MSLMs. Integrating insights from these aspects, we have developed a novel suite of models called Mipha. Compared to other MSLMs, our top-performing model, Mipha-3B, secures the highest performance in 9 out of 11 benchmarks. Moreover, when compared to the current state-of-the-art model, LLaVA-1.5-13B, Mipha-3B demonstrates comparable performance. We believe that Mipha offers a novel perspective on the training of strong MSLMs.

### 25 Mar 2024 v4

@@ Abstract
Multimodal Large Language Models (MLLMs) have showcased impressive skills in tasks related to visual understanding and reasoning. Yet, their widespread application faces obstacles due to the high computational demands during both the training and inference phases, restricting their use to a limited audience within the research and user communities. In this paper, we investigate the design aspects of Multimodal Small Language Models (MSLMs) and propose an efficient multimodal assistant named Mipha, which is designed to create synergy among various aspects: visual representation, language models, and optimization strategies. We show that without increasing the volume of training data, our Mipha-3B outperforms the state-of-the-art large MLLMs, especially LLaVA-1.5-13B, on multiple benchmarks. Through detailed discussion, we provide insights and guidelines for developing strong MSLMs that rival the capabilities of MLLMs. Our code is available at https://github.com/zhuyiche/llava-phi.
@@ 6 Conclusion
In this paper, we delve deeply into three aspects of Multimodal Small Language Models (MSLMs): the language models, visual representations, and optimization strategy. On the basis of our thorough analysis, we identify several improvements that can enhance the performance of MSLMs. Integrating insights from these aspects, we have developed a novel suite of models called Mipha. Compared to other MSLMs, our top-performing model, Mipha-3B, secures the highest performance in 8 out of 11 benchmarks. Moreover, when compared to the current state-of-the-art model, LLaVA-1.5-13B, Mipha-3B demonstrates comparable performance. We believe that Mipha offers a novel perspective on the training of strong MSLMs.
@@ C.1 More Experiments
C.1.3 More Experimental Result on Mipha.
Table 4: Multi-modal evaluation on instruction following benchmarks. POPE Li et al. [2023d]; MME-P: MME Perception Fu et al. [2023]; MME-C: MME Cognition Fu et al. [2023]; MMB: MMBench Liu et al. [2023d]; SEED: SEED-Bench Li et al. [2023b]; MM-Vet Yu et al. [2023]. ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
Method PT IT POPE MME- MMB SEED MM-Vet
P C en cn img
Multimodal Large Language Models
BLIP2-14B Li et al. [2023c] 129M - 85.3 1293.8 290.0 - - 49.7 22.4
InstructBLIP-8B Dai et al. [2023b] 129M 1.2M - - - 36 23.7 58.8 26.2
InstructBLIP-14B Dai et al. [2023b] 129M 1.2M 78.9 1212.8 291.8 - - - 25.6
Shikra-13B Chen et al. [2023b] 600K 5.5M - - - 58.8 - - -
IDEFICS-80B Laurençon et al. [2023] 353M 1M - - - 54.5 38.1 53.2 -
Qwen-VL Bai et al. [2023b] ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
Qwen-VL-Chat Bai et al. [2023b] ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
mPLUG-Owl2-7B Ye et al. [2023b] 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
LLaVA-7B Liu et al. [2023b] 595K 158K 72.9 809.6 247.9 38.7 36.4 37.0 25.5
LLaVA-1.5-7B Liu et al. [2023a] 558K 665K 85.9 1510.7 316.1 64.3 58.3 66.1 30.5
LLaVA-1.5-13B Liu et al. [2023a] 558K 665K 85.9 1531.3 295.4 67.7 63.6 68.2 35.4
Multimodal Small Language Models
MobileVLM-3B Chu et al. [2023] 558K 665K 84.9 1288.9 - 59.6 - - -
MobileVLM-v2-3B Chu et al. [2024] 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
MoE-LLaVA-3.6B Lin et al. [2024] 558K 1.59M 85.7 1431.3 - 68.0 - - 35.9
Bunny-3B He et al. [2024] 2M 695K 86.8 1488.8 289.3 68.6 - - -
Mipha-3B (Ours) 558K 665K 86.7 1488.9 295.0 69.7 64.3 68.9 32.1
[…]
Experimental Results on Instruction-Following Benchmarks. Furthermore, we evaluate our proposed Mipha on more challenging benchmarks that are designed for evaluation of MLLMs, i.e., POPE Li et al. [2023d], MME Fu et al. [2023], MMBench Liu et al. [2023d]. MMBench-CN Liu et al. [2023d], SEED-Bench-img Li et al. [2023b], and MM-Vet Yu et al. [2023]. As shown in Table LABEL:tbl:qa_table, our proposed model, Mipha-3B, exceeds LLaVA-1.5-13B in performance on MMBench and MMBench-CN by 2% and 0.7%, respectively, despite both models being trained on the same dataset. Additionally, Mipha-3B demonstrates stronger resistance to hallucinations, surpassing LLaVA-1.5-13B on POPE by 0.8%. In comparison with other MSLMs, Bunny-3B, MoE-LLaVA-3.6B Lin et al. [2024], and MobileVLM-v2-3B Chu et al. [2024], we are astonished to find that despite Mipha-3B utilizing the least amount of training data, it achieved the best performance in 4 out of 6 benchmarks. In particular, on MMBench, Mipha-3B outperformed Bunny-3B by 1.1% and MoE-LLaVA by 1.7%, which are combined with the same visual backbone SigLIP and LM Phi-2-2.7B respectively. These results validate the effectiveness of our optimization strategies.